# Dissecting the SWE-Bench Leaderboards: Profiling Submitters and Architectures of LLM- and Agent-Based Repair Systems

### v1

@@ 3.1.5. Open-Source Availability of Solutions
Open-source 44 22 28.16 47 38 16 39.5 65.8
Closed-source 24 18 35.34 55 40 22 47.8 68.2
@@
Table 6 presents the number of entries associated with open-source and closed-source solutions. In SWE-Bench Lite, the vast majority of entries (44 in total, representing 22 unique solutions) are open-source. In contrast, SWE-Bench Verified shows a reversed trend, with more closed-source entries (40) than open-source ones (28). As previously discussed, this leaderboard has attracted proportionally more submissions from industry actors, including startups and established companies, whose solutions are typically not released as open-source. We hypothesize that this tendency reflects a desire to protect the intellectual property of proprietary technologies.
The Fig.6(a) shows that closed-source submissions have consistently maintained the leading positions with few exceptions (before July 2024 and around November 2024). In SWE-Bench Verified the tend is different: As illustrated in Fig.6(b), throughout 2025, open-source submissions have constantly surpassed closed-source solutions in SWE-Bench Verified. However, the current state-of-the-art solutions is a close-source solution (NVIDIA Nemotron-CORTEXA)
Table 6 also reports the median and maximum % Resolved achieved by open- and closed-source solutions. In SWE-Bench Lite, both metrics are higher for closed-source entries compared to open-source ones. Figure 6(a) further illustrates this trend, showing that closed-source submissions have consistently held the leading positions, with only a few exceptions—namely, periods before July 2024 and around November 2024. In contrast, the trend in SWE-Bench Verified is different. As shown in Figure 6(b), In 2025, open-source and closed-source submissions have alternated in holding the top rank on SWE-Bench Verified. Nevertheless, the current state-of-the-art solution on SWE-Bench Verified remains a closed-source system (NVIDIA Nemotron-CORTEXA).
Answer to RQ 1 (Open & Closed source): While closed-source solutions exhibit a higher median % Resolved, emerging open-source tools have shown competitive performance, with several achieving state-of-the-art results on SWE-Bench Verified in 2025.
@@ 3.1.6. Large language Model Employed in Submissions
LLM Open Entries % Resolved Entries % Resolved
Total Median Max Total Median Max
Claude 3.5 Sonnet - 21 39 48.33 24 51.2 63.4
GPT-4o - 21 29.67 43 12 38.6 57.2
GPT-4 - 10 25.17 48.67 2 12.6 22.4
Qwen2.5 Y 3 23.67 24.67 9 30.2 63.4
Claude 3 - 4 18.5 26.33 2 12.6 18.2
LLama 3 Y 3 1.33 23.67 2 40.9 41.2
Claude Sonnet 3.7 - 0 - - 5 65.4 68.2
O3 Mini - 2 31.33 32.33 3 60.8 68.2
gpt-4o-mini - 0 - - 3 53.2 64.6
o1 - 0 - - 3 65.4 68.2
CodeLlama Y 0 - - 2 1.3 1.4
Deepseek Y 1 30.7 30.7 1 58.2 58.2
Gemini 2 Flash - 0 - - 2 48.2 52.2
Claude 2 - 1 3 3 1 4.4 4.4
GPT3/3.5 - 1 0.33 0.33 1 0.4 0.4
Granite Y 1 23.67 23.67 0 - -
Claude 3.5 Haiku - 0 - - 1 40.6 40.6
DeepSeek R1 Y 1 47 47 0 - -
DevStral Small Y 0 - - 1 46.8 46.8
o3 - 0 - - 1 68.2 68.2
Mistral-Large Y 1 23.67 23.67 0 - -
Table 7. Base LLMs used by the entries listed on SWE-Bench Lite and SWE-Bench Verified.
@@
SWE-Bench Lite SWE-Bench Verified
LLMs #E % Resolved LLMs #E % Resolved
Claude 3.5 Sonnet+DeepSeek R1 1 47 Claude Sonnet 3.7+o1+o3+O3 Mini 1 68.2
Claude 3.5 Sonnet 15 40.67 Claude Sonnet 3.7+o1 1 65.4
No Info 10 35.67 o1 1 64.6
Claude 3.5 Sonnet+GPT-4o 4 35 Claude 3.5 Sonnet+Qwen2.5 1 63.4
Claude 3.5 Sonnet+GPT-4 1 33 Claude Sonnet 3.7 3 63.2
O3 Mini 2 31.33 Claude 3.5 Sonnet+O3 Mini 1 60.8
Deepseek 1 30.7 gpt-4o-mini 2 58.9
GPT-4o 14 28.835 Claude 3.5 Sonnet+Deepseek 1 58.2
Claude 3+GPT-4o 2 25.83 Claude 3.5 Sonnet 15 51.6
GPT-4 8 25.17 Claude 3.5 Sonnet+gpt-4o-mini 1 48.6
Qwen2.5 2 24 Gemini 2 Flash 2 48.2
LLama 3+Mistral-Large+Qwen2.5+Granite 1 23.67 Claude 3.5 Sonnet+GPT-4o 5 47.2
GPT-4+GPT-4o 1 21.67 DevStral Small 1 46.8
Claude 3 2 8 No Info 18 45.3
Claude 2 1 3 O3 Mini 1 42.4
LLama 3 2 1.165 LLama 3 1 41.2
GPT3/3.5 1 0.33 Claude 3.5 Haiku 1 40.6
- - - LLama 3+Qwen2.5 1 40.6
- - - Qwen2.5 7 28.8
- - - GPT-4o 7 27
- - - Claude 3 2 12.6
- - - GPT-4 2 12.6
- - - Claude 2 1 4.4
- - - CodeLlama 2 1.3
- - - GPT3/3.5 1 0.4
Table 8. Combinations of base LLMs used per solution. #E is the number of entries that use the combination. % Resolved shows the median value.
Table 7 presents the base LLMs used by the submissions to both SWE-Bench Lite and SWE-Bench Verified. For each LLM, we also report summary statistics —median, maximum, and standard deviation— of the % Resolved achieved by the corresponding submissions. Figure 7 shows the distribution of the number of LLMs used per submission. Most entries from both leaderboards rely on a single LLM, while approximately ten use two. In addition, one entry in each leaderboard reports the use of four LLMs. A separate bar (“No Info”) summarizes entries for which the number of LLMs could not be determined. The combinations of LLMs used, their frequency and average % Resolved is shown in Table 8. Note that in this section we report the base LLM -pretrained, not fine-tuned used by each solution.
Performance of solutions per LLM
Let us first focus on Table 7. In both SWE-Bench Lite and SWE-Bench Verified, two models are predominantly used: Anthropic Claude 3.5 Sonnet (21 and 24 entries, respectively) and OpenAI GPT-4o (21 and 12 entries, respectively). In SWE-Bench Lite, state-of-the-art results are achieved using Claude 3.5 Sonnet (48.33%), GPT-4 (48.67%), and DeepSeek R1 (47%), the latter being used in only one submission in each leaderboard -in combination with other LLMs. In SWE-Bench Verified, the top results are obtained using Claude 3.7 Sonnet, OpenAI GPT-4o1 and OpenAI o3, all achieving 68.3%.
@@
While the majority of submissions report using a single LLM, as shown in Figure 7, Table 8 reveals that state-of-the-art results are often achieved through combinations of multiple LLMs. Notably, most LLM combinations appear in only one submission per leaderboard, with the exception of Claude 3.5 Sonnet + GPT-4o, which is used by 4 submissions in SWE-Bench Lite and 5 in SWE-Bench Verified.
@@
The majority of submissions rely on proprietary, closed LLMs accessed via API (see column Open). However, a variety of open-source models have also been employed. Examples include, Qwen2.5, the most used open-source LLM (employed by for instance SWE-Fixer), LLaMA, Mistral, and Granite (all of them used by IBM IA Agent), DeepSeek (used by Moatless Tool), and In general, the precision of approach using only open-source LLMs is lower than those using closed LLMs. There are, nonetheless, exceptions such as DeepSeek used by Moatless which achieves 47% on Lite, a results that is near to state-of-the-art (48%).
On SWE-Bench Verified, the performance of approaches based solely on open-source LLMs does not reach state-of-the-art levels. We observe that these models are rarely used alone; instead, they are often combined with closed-source LLMs. For instance, a near state-of-the-art submission —AgentScope, with 63.4% precision— employs a combination of two models: Claude 3.5 Sonnet, used in various stages such as code modification, and Qwen 2.5, leveraged by the reward model to select the best outcome among multiple trials.
@@
One of the key advantages of using open-source LLMs is the ability to fine-tune them for specific tasks. Several submissions leveraged this flexibility to tailor models for different components of the repair pipeline. For example, SWE-Fixer (xie2025SWEFixerTrainingopensourcellms, ) fine-tuned Qwen 2.5 twice: once for the bug retriever and once for the patch editor. Nebius fine-tuned two separate models: Qwen 2.5 for patch generation and LLaMA 3.1 for the patch critic. DARS (aggarwal2025darsdynamicactionresampling, ) fine-tuned DeepSeek R1 to serve as a patch reviewer, while AgentScope fine-tuned Qwen-2.5-Coder specifically for patch assessment. SWE-agent-LM-32B applied fine-tuning to Qwen 2.5 using the SWE-smith dataset (yang2025swesmith, ), which was generated from trajectories of SWE-agent (yang2024sweagentagent, ) running Claude 3.7 Sonnet. DevStral is a fine-tuned version of Mistral-Small-3.1, trained to address real-world GitHub issues. Lingma SWE-GPT (ma2024lingmaswegpt, ) fine-tuned Qwen 2.5 for its specialized software engineering tasks.
Meta presented the approach SWE-RL (wei2025SWE-RL, ) for enabling LLMs to autonomously recover a developer’s reasoning processes and solutions. The resulting reasoning model, Llama3-SWE-RL-70B fine-tuned Llama 3 using Reinforcement learning (RL). During the training, the policy LLM is tasked with solving a given issue through reasoning and producing the code changes, then converted into a consistent patch format for reward calculation. SWE-RL helps the LLM develop reasoning skills like self-reflection, exploring alternatives, and divide-and-conquer strategies.
Some of these fine-tuned models, such as DevStral, Lingma SWE-GPT, SWE-agent-LM-32B, there are publicly available on model repositories such as HuggingFace or ModelScope. Others, such the case of DARS or the SWE-Fixer Editor/Retriever models, in addition of publishing the models on HuggingFace, also uploaded the training data on this platform [^17] https://huggingface.co/AGENTDARS, https://huggingface.co/datasets/SWE-bench/SWE-smith-trajectories.
@@
We detect 8 entries fully based on open-source LLMs, including: 1) SWE-Agent + SWE-agent-LM-32B; 2) SWE-Fixer; 3) Nebius AI; 4) IBM AI Agent SWE-1.0, which report to use four different LLMs including their LLM IBM Granite, the only entry that use it; 5) Open-Hands + DevStral Small; 6) Moatless Tools + Deepseek V3; and 7) Lingma SWE-GPT; 8) SWE-RL; As mentioned earlier, open-source models allow for fine-tuning, enabling task-specific customization.
Beyond the openness of the LLMs and the public availability of fine-tuned models, some submitters have also made their full solution code publicly available (e.g., agent scaffolds). Some of these submissions correspond to SWE-Agent, Lingma Agent, SWE-RL and SWE-Fixer.
This level of openness means that, given appropriate infrastructure to deploy a LLM, others can fully replicate and build upon these systems, promoting transparency, reproducibility, and further research.
@@
Answer to RQ 1 (LLMs): Proprietary models such as Claude 3.5/3.7 Sonnet and GPT-4o/4 currently lead in achieving state-of-the-art results on SWE-Bench. Solutions that combine closed and open-source LLMs—including those that incorporate fine-tuned models—also demonstrate competitive performance. Approaches based solely on fine-tuned open-source LLMs tend to yield more modest precision but offer benefits in terms of transparency, reproducibility and adaptability.
@@
3.2. RQ: What is the high-level architecture of the approaches behind the submissions to the SWE-Bench Lite and SWE-Bench Verified leaderboards?
In this section, we classify the approaches according to three key dimensions: 1) workflow authoring, 2) control flow autonomy, and 3) the total number of agents that comprise the approach.
@@
Rather than examining each dimension separately, we explore how these dimensions are combined within each solution’s design This leads to the identification of seven architectural groups (G1-G7), along with an additional group (G8) for submissions lacking sufficient information for classification . For each approach, we report the performance as listed on the leaderboards (column % Resolved). In cases where multiple submissions from a same approach exist (such as those using different LLMs) we just report the highest score.

### v2

@@ 3.1.4. Open-Source Availability of Solutions
Open-source 53 27 30.33 60.33 52 23 44.2 75.2
Closed-source 26 20 35.67 55 47 24 50 73.2
@@
Table 4 presents the number of entries associated with open-source and closed-source solutions. In SWE-Bench Lite, the 67% of entries (53 in total, representing 27 unique solutions) are open-source. In contrast, SWE-Bench Verified that percentage is lower: 52%. As previously discussed, this leaderboard has attracted proportionally more submissions from industry actors, including startups and established companies, whose solutions are typically not released as open-source. In both leaderboards, entries associated with closed-source systems exhibit a higher median % Resolved. However, the top-ranked entries in both SWE-Bench Lite and SWE-Bench Verified are open-source.
Temporal Evolution of % Resolved
Fig.7 shows the evolution of open- and closed- source entries over time. In SWE-Bench Lite, until early 2025, state-of-the-art results were achieved exclusively by closed-source solutions. However, since then, just few close-solution were submitted and open-source solution recently achieve state-of-the-art results. In SWE-Bench Verified the trend is different: As illustrated in Fig.7(b), throughout 2025, open-source and closed-source submissions have alternated in holding the top rank.
Answer to RQ 1 (Open & Closed source): While closed-source solutions exhibit a higher median % Resolved, emerging open-source tools are showing competitive performance, with several of them achieving state-of-the-art results on both SWE-Bench Lite and SWE-Bench Verified in 2025.
@@ 3.1.5. Large language Model Employed in Submissions
LLMs #E % Resolved LLMs #E % Resolved
Med Max Med Max
Claude 4 Sonnet 2 58.5 60.33 Claude 3.7 Sonnet+Claude 4 Sonnet+Claude 4 Opus+Gemini 2.5 Pro 1 75.2 75.2
Claude 3.7 Sonnet+o4-mini 1 60 60 Claude 4 Sonnet+o4-mini 1 74.4 74.4
No Info 9 35.67 55 Claude 4 Opus 1 73.2 73.2
Claude 3.7 Sonnet+Gemini 2.5 Pro 1 51.67 51.67 Claude 4 Sonnet 5 70.4 72.4
Claude 3.5 Haiku+Claude 3.5 Sonnet+Gemini 2.5 Pro 1 49 49 Claude 3.7 Sonnet+Claude 4 Sonnet+GPT-4+Gemini 2.5 Pro 1 71 71
GPT-4 8 25.17 48.67 Claude 3.7 Sonnet+o1+o4-mini+Gemini 2.5 Pro 1 70.6 70.6
Claude 3.5 Sonnet 17 41 48.33 Claude 3.7 Sonnet+o3+o4-mini 1 70.4 70.4
Claude 3.5 Sonnet+o3-mini+o4-mini 1 48.33 48.33 Claude 3.7 Sonnet+o3+Gemini 2.5 Pro 1 70.2 70.2
Claude 3.7 Sonnet 1 48 48 Claude 3.7 Sonnet+o4-mini 1 70 70
Claude 3.5 Sonnet+DeepSeek R1 1 47 47 Claude 3.7 Sonnet+o1+o3+o3-mini 1 68.2 68.2
Claude 3.5 Sonnet+GPT-4o 4 35 43 Claude 3.7 Sonnet 3 63.2 66.4
Claude 3.5 Haiku+Claude 3.5 Sonnet 1 41.67 41.67 No Info 18 45.3 65.8
GPT-4o 14 28.835 39.33 Claude 3.7 Sonnet+o1 1 65.4 65.4
Deepseek V3 (*) 2 33.685 36.67 o1 1 64.6 64.6
Claude 3.5 Sonnet+GPT-4 1 33 33 gpt-4o-mini 2 58.9 64.6
o3-mini 2 31.33 32.33 Claude 3.5 Sonnet+Qwen2.5 1 63.4 63.4
Claude 3+GPT-4o 2 25.83 26.33 Claude 3.5 Sonnet 15 51.6 62.8
Qwen2.5 (*) 3 23.33 24.67 Claude 3.5 Sonnet+o3-mini 1 60.8 60.8
LLama 3+Mistral-Large+Qwen2.5+Granite (*) 1 23.67 23.67 Qwen3 (*) 2 50.5 58.8
GPT-4+GPT-4o 1 21.67 21.67 Claude 3.5 Sonnet+Deepseek 1 58.2 58.2
Claude 3 2 8 11.67 Claude 3.5 Sonnet+GPT-4o 5 47.2 57.2
Claude 2 1 3 3 Gemini 2 Flash 2 48.2 52.2
LLama 3 (*) 2 1.165 1.33 Claude 3.5 Sonnet+gpt-4o-mini 1 48.6 48.6
GPT3/3.5 1 0.33 0.33 Qwen2.5 (*) 11 30.2 47
- - - - DevStral Small (*) 1 46.8 46.8
- - - - o3-mini 1 42.4 42.4
- - - - Amazon Nova Premier 1 42.4 42.4
- - - - LLama 3 (*) 1 41.2 41.2
- - - - Claude 3.5 Haiku 1 40.6 40.6
- - - - LLama 3+Qwen2.5 (*) 1 40.6 40.6
- - - - GPT-4o 7 27 38.8
- - - - GPT-4 2 12.6 22.4
- - - - Claude 3 2 12.6 18.2
- - - - Claude 2 1 4.4 4.4
- - - - CodeLlama (*) 2 1.3 1.4
- - - - GPT3/3.5 1 0.4 0.4
Table 5. Combinations of base LLMs used per solution. #E is the number of entries that use the combination. % Resolved shows the median and maximum values. An asterisk (*) indicates that all models are open-source.
@@
Table 5 presents the base LLMs used in the submissions to both SWE-Bench Lite and SWE-Bench Verified, and how frequent these LLMs are used together in the submissions. The table is sorted by % Resolved. As usual, some entries did not include enough information (see separate bar “No Info” in the figure.
In SWE-Bench Lite, Claude 4 Sonnet employed by the approach ExpeRepair (mu2025experepairdualmemoryenhancedllmbased, ) achieved the highest performance with a preposition of 60.3%. This model has been used in two entries in total (column #E): it was also used by SWE-agent. Claude 3.5 Sonnet is the most frequently used LLM, appearing in 16 entries as a standalone model, and in several others in combination with additional LLMs. For instance, four entries combine both Claude 3.5 Sonnet and GPT-4o.
In SWE-Bench Verified, a similar trend is observed: the state-of-the-art result (75.2%), by TRAE (June 2025) was achieved using a combination of multiples state-of-the-art models including Claude 4 Sonnet and Opus Claude 3.7 Sonnet. Nevertheless, using Claude 4 Sonnet or Claude 4 Opus individually also led to competitive performance. Claude 3.5 Sonnet remains the most frequently used model, appearing in 17 entries as a standalone LLM and in five entries combined with others, such as GPT-4o. The second most frequently used LLM in isolation is Qwen2.5, with 11 entries. Notably, it is an open-source model; however, the solutions employing it achieved lower performance, with precision scores up to 47%.
@@
The majority of submissions report using a single LLM, as shown in Figure 8. Excluding the most recent state-of-the-art models from the Claude 4 family, Table 5 shows that top-performing results are often achieved through combinations of multiple LLMs. Notably, most LLM combinations appear in only one submission per leaderboard, with a few exceptions: Claude 3.5 Sonnet + GPT-4o is used by four submissions in SWE-Bench Lite and five in SWE-Bench Verified.
Solutions that use more than one LLM often assign different roles to each model within the repair process. For example, the second-best solution in SWE-Bench Lite, Replit.ai Agent, employs Claude 3.7 Sonnet for orchestration and decision-making, and o4-mini for deeper analysis and code generation. Similarly, InfantAgent (lei2024InfantAgent, ) leverages closed-source models exclusively as the “brain” of its agent-based architecture, while smaller open-source models are employed for simpler, repetitive tasks.
Another example is Devlo, which achieved a competitive 70.2% on SWE-Bench Verified by using three LLMs —Claude 3.7 Sonnet, o3, and Gemini 2.5 Pro— during the patch generation phase to promote diversity in candidate patches. In some cases, while one or more LLMs are responsible for generating patches, a different LLM is used for analyzing and selecting among them —a strategy referred to as “LLM-as-a-judge” or “LLM-as-a-selector”, respectively. For instance, a submission of TRAE (which achieved 70.4% in May 2025) uses o1 to select among patches generated by three different models: Claude 3.7 Sonnet, Gemini 2.5 Pro and o4-mini. Similarly, AgentScope, with 63.4% precision— uses the open-source Qwen2.5 to select the best result among multiple trials made with Claude 3.5 Sonnet.
@@
The majority of submissions rely on proprietary, closed-source LLMs accessed via API, most notably from the Claude 3 and 4 model families. However, a range of open-source models has also been employed, typically after fine-tuning for specific tasks such as patch evaluation. Combinations of fully open-source models are indicated in Table 5 with an asterisk (*). The most frequently used open-source model is Qwen2.5, serving as the sole model in 3 entries from SWE-Bench Lite and 11 from SWE-Bench Verified. Some entries rely exclusively on a set of open-source models, for example, one submission from IBM uses LLama, Mistral-Large, Qwen2.5 and Granite. In general, the precision of approach using only open-source LLMs is lower than those using closed LLMs, in both leaderboards. Nevertheless, some solutions combine both closed- and open-source models. One such example is AgentScope, which achieved near state-of-the-art performance on SWE-Bench Verified (63.4%), at the time the result was published (February 2025), using Claude 3.5 Sonnet in combination with Qwen2.5. Notably, no other submission using Claude 3.5 Sonnet outperformed AgentScope, suggesting that properly fine-tuned open-source models can enhance the effectiveness of solutions built around closed-source LLMs.
@@
One of the key advantages of using open-source LLMs is the ability to fine-tune them for specific tasks. Several submissions leveraged this flexibility to tailor models for different components of the repair pipeline. For example, SWE-Fixer (xie2025SWEFixerTrainingopensourcellms, ) fine-tuned Qwen 2.5 twice: once for the bug retriever and once for the patch editor. Nebius fine-tuned two separate models: Qwen 2.5 for patch generation and LLaMA 3.1 for the patch critic. DARS (aggarwal2025darsdynamicactionresampling, ) fine-tuned DeepSeek R1 to serve as a patch reviewer, while AgentScope, as mentioned before, fine-tuned Qwen-2.5 specifically for patch assessment. SWE-agent-LM-32B applied fine-tuning to Qwen 2.5 using the SWE-smith dataset (yang2025swesmith, ), which was generated from trajectories of SWE-agent (yang2024sweagentagent, ) running Claude 3.7 Sonnet. DevStral is a fine-tuned version of Mistral-Small-3.1, trained to address real-world GitHub issues. Lingma SWE-GPT (ma2024lingmaswegpt, ) fine-tuned Qwen 2.5 for its specialized software engineering tasks.
Meta presented the SWE-RL approach (wei2025SWE-RL, ) for enabling LLMs to autonomously recover a developer’s reasoning processes and solutions. The resulting reasoning model, Llama3-SWE-RL-70B, fine-tuned Llama 3 using Reinforcement learning (RL). During the training, the policy LLM is tasked with solving a given issue through reasoning and producing the code changes, then converted into a consistent patch format for reward calculation. SWE-RL helps the LLM develop reasoning skills like self-reflection, exploring alternatives, and divide-and-conquer strategies. Co-PatcheR (tang2025copatchercollaborativesoftwarepatching, ) fine-tuned three distinct instances of Qwen2.5, each tailored for a specific task —localization and generation, and test generation with and without assertions— using distillation data generated by Claude 3.7 Sonnet with explicit reasoning chains.
Some of these fine-tuned models, such as DevStral, Lingma SWE-GPT, SWE-agent-LM-32B, there are publicly available on model repositories such as HuggingFace or ModelScope. Others, such the case of the approaches DARS and SWE-Fixer, also uploaded the training data on this platform [^14] https://huggingface.co/AGENTDARS, https://huggingface.co/datasets/SWE-bench/SWE-smith-trajectories.
@@
Beyond the openness of the LLMs and the public availability of fine-tuned models, some submitters have also made their full solution code publicly available (e.g., agent scaffolds). The open-source approaches, with at least one submission exclusively using open-source models, are 13: 1) AgentScope, 2) Alibaba Lingma Agent, 3) Co-PatcheR, 4) DARS Agent, 5) DeepSWE, 6) KGCompass, 7) MCTS-Refine, 8) Moatless Tools, 9) OpenHands, 10) RAG, 11) SWE-Fixer, 12) SWE-agent, and 13) Skywork-SWE. This level of openness means that, given appropriate infrastructure to deploy a LLM, others can fully replicate and build upon these systems, promoting transparency, reproducibility, and further research.
Temporal Evolution of % Resolved per LLM employed
Figure 9 shows the precision over time for each set of LLMs used together, based on their publication dates on the leaderboards. For clarity, Figure 9(a) displays the top 70 out of 79 entries from SWE-Bench Lite, while Figure 9(b) shows the top 45 out of 99 entries from SWE-Bench Verified. The colors represent: (1) Blue: submissions just using Anthropic Claude models, (2) Green: submissions just using OpenAI models, (3) Purple: submissions using Anthropic and OpenAI, but not others, (4) Yellow: submissions only using open-source models, (5) Orange: other cases such as submissions using other vendors, or combinations between closed- and open-source models, and finally, (6) Black: submissions without specifying the model used.
[…]
In SWE-Bench Verified, until January 2025, state-of-the-art submissions either used Claude 3.5 Sonnet (++) or did not disclose the underlying models (X). In January, the submission from W&B ($B$) reached the top precision using o1 (64.6%), holding the lead until the first submission combining Claude 3.7 Sonnet and o1 ($A$). Since May 2025, all leading submissions on the leaderboard have included models from the Claude 4 series. The first of these was a submission by Anthropic using only Claude 4 Opus ($D$), which achieved 73.20%. Subsequent top submissions combined Claude 4 with other LLMs ($B$), including the current leader by TRAE, which reached 75.2%
@@
Answer to RQ 1 (LLMs): Approaches leveraging proprietary models, first Claude 3.5 Sonnet and more recently Claude 4 variants, have consistently achieved the highest precision on both leaderboards. Solutions that combine closed and open-source LLMs—including those that incorporate fine-tuned models—also demonstrate competitive performance. Approaches based solely on fine-tuned open-source LLMs tend to yield more modest precision but offer benefits in terms of transparency, reproducibility and adaptability.
(a) SWE-Bench Lite
(b) SWE-Bench Verified
Figure 9. Evolution of % Resolved by LLMs used.
@@
3.2. RQ2: What is the high-level architecture of the approaches behind the submissions to the SWE-Bench Lite and SWE-Bench Verified leaderboards?
@@
In this section, we classify the approaches according to three key dimensions that we presented in detail in 2.2: 1) workflow authoring, 2) control flow autonomy, and 3) the total number of agents that comprise the approach. Rather than examining each dimension separately, we explore how these dimensions are combined within each solution’s design. This leads to the identification of seven architectural groups (G1-G7) [^15] Some combinations are not possible, such as an agent performing a fixed execution., along with an additional group (G8) for submissions lacking sufficient information for classification. These groups are:
  • G1: Human-Workflow with Fixed execution - No Agent (Section 3.2.1).
  • G2: Human-Workflow with Fixed execution - with Single Agent (Section 3.2.2).
  • G3: Human-Workflow with Fixed execution - with Multiple Agents (Section 3.2.3).
  • G4: Human-Workflow with Scaffolded Execution - Single Agent (Section 3.2.4).
  • G5: Human-Workflow with Scaffolded Execution - Multiple Agents (Section 3.2.5).
  • G6: Workflow Emergent and Emergent Autonomy - Single Agent (Section 3.2.6),
  • G7: Workflow Emergent and Emergent Autonomy - Multiple Agents (Section 3.2.7).
  • G8: Unclassified Submissions (Section 3.2.8).
For each approach, we enumerate the different approaches, reporting their performance as listed in the leaderboards (column % Resolved). In cases where multiple submissions from the same approach exist (such as those using different LLMs), we just report the highest score. Finally, in Section 3.2.9, we analyze the performance of each of these groups.